OpenStudy (anonymous):

solve by the elimination method, type an ordered pair 7r-8s=55 8r+7s=79

OpenStudy (anonymous):

You need to get either the r's or the s's to be eliminated. Because the s's have a "+" and a "-" already, I would choose the s's.

OpenStudy (anonymous):

The first common number 7 and 8 have in common is 56 so we are going to multiply the first equation by 7 so I can get -56s

OpenStudy (anonymous):

Going to multiply the second equation by 8 so we can get a +56s

OpenStudy (anonymous):

7(7r - 8s = 55) 8(8r + 7s = 79)

OpenStudy (anonymous):

49r - 56s = 385 64r + 56s = 632

OpenStudy (anonymous):

i solve that quastion yesterday Charlie... http://openstudy.com/?F240238930417BYII5Z=_#/updates/4eab39f4e4b02eee39e6ea69

OpenStudy (anonymous):

Now add these up 113r = 1017 r = 9

OpenStudy (anonymous):

Yes, Leo you were very helpful yesterday and I appreciate it. Only I discovered I also need a set of ordered pairs out of the equation. Thanks dude.

OpenStudy (anonymous):

Substitue 9 in for r 7(9) -8s = 55 63 - 8s = 55 -8s = -8 s = 1 Ordered pair (9,1)

OpenStudy (anonymous):

thank you very much blexting.

OpenStudy (anonymous):

was just put the answer in the form (r, s)

OpenStudy (anonymous):

:)